OpenStudy (truwhovian11):

medal+fan for the first answer from the set(7,10,13) use substitution to determine which value of X makes the equation true. 80-4x=40 please explain A.10 B.13 C.none D.7

OpenStudy (truwhovian11):

@TheCalcHater @ace-n-it

OpenStudy (thecalchater):

Ok so plug in you values a b c d and see what u get.

OpenStudy (truwhovian11):

excuse me what

OpenStudy (mathmale):

"from the set(7,10,13) use substitution to determine which value of X makes the equation true. 80-4x=40" For example, choose x=13. This is one of the given set of input values. Substitute 13 for x in the equation above. Is the equation then true or false? What would your conclusion about x=13 be?

OpenStudy (truwhovian11):

um idk ;-;

OpenStudy (thecalchater):

where x is replace that with one of your values and compute it.

OpenStudy (truwhovian11):

can you please eplain

OpenStudy (truwhovian11):

@TheCalcHater

OpenStudy (thecalchater):

ok so when we substitute for x we are just plugging in so 80-4(x)=40 so put in a for x so 80-4(10)=40 is that true?

OpenStudy (truwhovian11):

yes